OpenStudy (anonymous):

if f(x)=-6+12, then f^-1(x)= ?

OpenStudy (anonymous):

I assume you mean f(x) = -6x + 12 To find the inverse, intercjange x with y and y with x AND then solve for y. We are given f(x) = y = -6x + 12 Inverse will be: x = -6y + 12 Solve for y, you get y = (x-12)/(-6) we can say that the inverse of f, denoted by f^-1(x) = (x-12)/(-6)

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

welcome.